# Sparse identification of quasipotentials via a combined data-driven method

### v1

@@ Abstract
The quasipotential function allows for comprehension and prediction of the escape mechanisms from metastable states in nonlinear dynamical systems. This function acts as a natural extension of the potential function for non-gradient systems and it unveils important properties such as the maximum likelihood transition paths, transition rates and expected exit times of the system. Here, we leverage on machine learning via the combination of two data-driven techniques, namely a neural network and a sparse regression algorithm, to obtain symbolic expressions of quasipotential functions. The key idea is first to determine an orthogonal decomposition of the vector field that governs the underlying dynamics using neural networks, then to interpret symbolically the downhill and circulatory components of the decomposition. These functions are regressed simultaneously with the addition of mathematical constraints. We show that our approach discovers a parsimonious quasipotential equation for an archetypal model with a known exact quasipotential and for the dynamics of a nanomechanical resonator. The analytical forms deliver direct access to the stability of the metastable states and predict rare events with significant computational advantages. Our data-driven approach is of interest for a wide range of applications in which to assess the fluctuating dynamics.
[…]
@@ 1 Introduction
Examining the stability of attractors and investigating noise-induced switching between metastable solutions is indispensable for understanding significant biophysical, engineering and ecological phenomena [1, 2], including the folding of proteins [3], the kinetics of chemical reactions [4], and the dynamics of mechanical devices [5, 6]. Comprehensive knowledge of activated escape from a metastable state is crucial for mastering non-equilibrium phenomena that arise from the interplay between nonlinearity and fluctuations [7, 8, 9].
@@
An effective method for illustrating the physics of fluctuation-activated transitions between coexisting stable states is the so-called “ball-in-cup” diagram analogy. In this representation, the state of the dynamical system is depicted by the ball’s position as it rolls across a surface. Random perturbations have the potential to shift the ball from one basin of attraction to another. This surface is defined by a potential function, serving as a natural metric for stability. However, it is worth noting that systems suitable for this depiction, known as gradient systems, are relatively rare [10]. Indeed, when attempting to model the non-equilibrium statistics of transitions between stable states, it is common to encounter non-gradient systems [11]. In this case, the transition dynamics becomes predictable and is entirely defined by the quasipotential that extends the idea of energy functions [12]. As a result, the quasipotential landscape provides an intuitive picture of the fundamental dynamical characteristics of complex systems operating outside of equilibrium [13]. With the quasipotential at hand, valuable asymptotic information can be easily estimated. For instance, a calculation over the energy surface allows for the straightforward determination of the maximum likelihood transition path [14]. Additionally, a precise estimate for the invariant probability distribution near an attractor and the expected escape time from a basin of attraction has been formulated on the basis of the quasipotential topology [15].
Although the quasipotential holds considerable importance for being the informative landscapes in stochastic dynamical systems, calculating it presents a formidable challenge. As a matter of fact, it can be analytically derived only in in some special cases [16]. Ono can rely on various minimum action methods designed to compute the quasipotential between two specified points by solving a path-space minimization problem [17]. Differently, others focus on solving a functional minimization problem, aimed to assess the quasipotential landscapes placing emphasis on the action functional. Nevertheless, all the approaches, such as the variational formulation based on the Freidlin-Wentzell action functional or solving the associated Hamilton-Jacobi equation over a mesh of spatial points, pose difficult numerical challenges with prohibitive computational expense. This is where machine learning (ML) techniques enter overwhelming into the picture. Even if still in early stage of development, data-driven methods and ML-based approaches have significantly impacted the realm of nonlinear dynamics. The emerging scaling of ML techniques have demonstrated capabilities in both interpretation and development of new identification methods [18, 19, 20].
@@
The quasipotential is identified via ML in Lin et al. [21], where an efficient neural network approach is employed to find the orthogonal decomposition of the vector field that governs the dynamics. The idea of utilizing a decomposition of the vector field was also used to compute the generalized potential or invariant distribution for randomly perturbed dynamical systems [22, 23]. In [24], a neural network with automatic differentiation inspired by physics-informed neural network is employed to compute the quasipotential based on the Hamilton-Jacobi equation. A neural network estimation of the quasipotential also promotes for the strategic control of the mean exit time for general stochastic dynamical systems [25].
In this study, we introduce a new data-driven approach to achieve a symbolic interpretation of the quasipotential function. The neural network-based approach of Lin et al. and the sparse regression identification technique developed by [26], are combined to find an analytical model. We leverage the capabilities of neural networks to find the orthogonal decomposition of the vector field driving the dynamics. Subsequently, the functions are reinterpreted using constrained regression within a suitable basis of functions. The quasipotential equation has parsimonious and interpretable form, in the sense that the model is sparse in the space of all the possible functions [27, 28]. In our method, observed trajectories of the dynamical system are the only input required by the combined approach and the goal is to identify a symbolic quasipotential function using a fully data-driven methodology.
The paper is organized as follows. In Sec. 2, we introduce the background of the quasipotential for non-gradient systems and the two parts of the data-driven method. In Sec. 3, two applications of the method are presented. In Sec. 4, we draw the conclusions.
@@ 2.1 Expanding on the concept of potential function
We undergo the investigation and interpretation of the process through which a dynamical systems evolve when in presence of multiple equilibra. The underlying physics of transient dynamics and attractors is assumed to be described by the state equation
@@
$\mathbf{f}=-\nabla U(\mathbf{x})$ (2)
we are in the presence of a gradient systems, with $U$ referred to as the “potential”. Thanks to this function, the dynamics of this system can be likened to that of a ball-in-cup setup, where the surface is defined by the potential. In a more general scenario, i.e. non-gradient systems, it is not possible to find $U$ such that Eq. (2) is satisfied. However, the concept of the potential function can be extended to encompass non-gradient systems via the Freidlin-Wentzell quasipotential function [12]. To determine this function, we decompose the vector field $\mathbf{f}$ as
$\mathbf{f}(\mathbf{x})=-\nabla V(\mathbf{x})+\mathbf{g}(\mathbf{x})$ (3)
@@
Under suitable extra conditions for $V$ [12, 21], the function $V$ is locally a scalar multiple of a function known as the Freidlin-Wentzell quasipotential $U$, i.e. $U=2V$. It represents a useful generalization of the potential function for a n-dimensional non-gradient system. In terms of the vector field, the $-\nabla V$ component in Eq. (3) spurs a ball to roll to the bottom of a valley in the ball-in-a-cup analogy, thus $V$ specifies an hyper-surface in which all the trajectories move “downhill” in the absence of perturbations. From Eq. (3) and the Hamilton-Jacobi relation (4) it is straightforward to verify that
@@
hence, $\mathbf{g}$ and $\nabla V$ are perpendicular. Without additional external forces, the “circulatory” component $\mathbf{g}$ creates the circulation of trajectory around levels of $V$. For convenience, we refer to $\nabla V(\mathbf{x})$ and $\mathbf{g}(\mathbf{x})$ in the decomposition (3), (5) as downhill and circulatory components, respectively.
Our method derives $V$ without solving numerically Eq. (4). A two-step data-driven approach is designed to output the symbolic expression of the quasipotential in terms of the state space variables, i.e. $V(\mathbf{x})$. First, a neural network determines the elements in the orthogonal decomposition of Eq. (3) (Sec. 2.2), second, the learned mapping functions are reinterpreted via a constrained regression (Sec. 2.3).
@@ 2.2 Dataset and neural networks
$L(\theta)=$ $\frac{1}{NM}\sum_{i,j}\bigg{\lVert}\frac{1}{h}\Big{(}\mathcal{I}_{h}(X_{i}^{j,0};\mathbf{f}_{\theta})-X_{i}^{j,h}\Big{)}\bigg{\rVert}_{2}^{2}+\frac{\lambda}{S}\sum_{k=1}^{S}w\left(\frac{\nabla V_{\theta}(\tilde{X}_{k})\cdot\mathbf{g}_{\theta}(\tilde{X}_{k})}{|\nabla V_{\theta}(\tilde{X}_{k})|\cdot|\mathbf{g}_{\theta}(\tilde{X}_{k})|},\delta\right),$ (7)
@@
in which the piecewise function $w(z,\delta)=z^{2}\mathbb{I}_{z\geq 0}+\delta z^{2}\mathbb{I}_{z<0}$ with $\delta=0.1$, $\lambda$ is a parameter controlling the strength of the loss term corresponding to the orthogonality condition, and $\{\tilde{X}_{k}\}$, $1\leq k\leq S$ is a representative subset of $\{X_{i}^{j,0}\}$ with a uniform distribution in the state space [21].
@@
We parameterize the potential component $V(\mathbf{x})$ by a sum of a vanilla neural network $V^{\text{NN}}_{\theta}(\mathbf{x})$ with hyperbolic tangent ($\tanh$) as the activation function and a harmonic function, to make the learned system preserve good physical properties [21]. In addition, the training performances with the loss function (7) are ameliorated by adding: i) a normalization step to the input layer of $V_{\theta}$, $\mathbf{g}_{\theta}$ with respect to the data points; ii) a scalar multiplication to the output of the networks. In details, for the network $V_{\theta}$ we have
@@
$V_{\theta}(\mathbf{x})$ $=\eta_{v}\cdot\tilde{V}_{\theta}\left(\frac{\mathbf{x}-\mu}{\sigma}\right),$ (8)
$\tilde{V}_{\theta}(\mathbf{y})$ $=V^{\text{NN}}_{\theta}(\mathbf{y})+|\mathbf{y}|^{2},$
with $\mu$ and $\sigma$ being the sample mean and deviation of the data points $\{X_{i}^{j,0}\}$, respectively. The parameter $\eta_{v}$ rescales the magnitude of $|\nabla V_{\theta}|$ to that of $|\mathbf{f}|$ estimated from the data. This is obtained by
$\eta_{v}$ $=\operatorname*{arg\,min}_{\eta>0}\sum_{i,j}\Big{(}\big{|}\eta\mathbf{v}_{i}^{j}\big{|}-\big{|}Y_{i}^{j}\big{|}\Big{)}^{2}$ (9)
$=\sum_{i,j}\big{|}\mathbf{v}_{i}^{j}\big{|}\cdot\big{|}Y_{i}^{j}\big{|}\bigg{/}\sum_{i,j}\big{|}\mathbf{v}_{i}^{j}\big{|}^{2},$
where the vector $\mathbf{v}_{i}^{j}=\nabla\tilde{V}_{\theta}\big{(}(X_{i}^{j,0}-\mu)/\sigma\big{)}$ and $Y^{j}_{i}=\big{(}X_{i}^{j,h}-X_{i}^{j,0}\big{)}/h$ is the estimated vector for $\mathbf{f}$ at the state $X_{i}^{j,0}$. Similarly, by estimating a multiplying scalar $\eta_{g}$ such that the magnitude of $|\mathbf{g}_{\theta}|$ approximates that of $|\mathbf{f}|$, the network $\mathbf{g}_{\theta}$ is
@@
$\mathbf{g}_{\theta}(\mathbf{x})=\eta_{g}\cdot\mathbf{g}^{\text{NN}}_{\theta}\left(\frac{\mathbf{x}-\mu}{\sigma}\right).$ (10)
Note that the parameters $\mu$, $\sigma$, $\eta_{v}$ and $\eta_{g}$ are kept as constant during the entire training process.

### v2

@@ Abstract
The quasipotential function allows for comprehension and prediction of the escape mechanisms from metastable states in nonlinear dynamical systems. This function acts as a natural extension of the potential function for non-gradient systems and it unveils important properties such as the maximum likelihood transition paths, transition rates and expected exit times of the system. Here, we demonstrate how to discover parsimonious equations for the quasipotential directly from data. Leveraging machine learning, we combine two existing data-driven techniques, namely a neural network and a sparse regression algorithm, specifically designed to symbolically describe multistable energy landscapes. First, we employ a vanilla neural network enhanced with a renormalization and rescaling procedure to achieve an orthogonal decomposition of the vector field. Next, we apply symbolic regression to extract the downhill and circulatory components of the decomposition, ensuring consistency with the underlying dynamics. This symbolic reconstruction involves a simultaneous regression that imposes constraints on both the orthogonality condition and the vector field. We implement and benchmark our approach using an archetypal model with a known exact quasipotential, as well as a nanomechanical resonator system. We further demonstrate its applicability to noisy data and to a four-dimensional system. Our model-unbiased analytical forms of the quasipotential is of interest to a wide range of applications aimed at assessing metastability and energy landscapes, serving to parametrically capture the distinctive fingerprint of the fluctuating dynamics.
keywords:
Quasipotential, data-driven identification, neural-networks, sparse regression
@@ 1 Introduction
Examining the stability of attractors and investigating noise-induced switching between metastable solutions is indispensable for understanding critical phenomena in biophysics, engineering, and ecology [Beisner2003, Nolting2016]. These include processes such as protein folding [wales2003energy], the kinetics of chemical reactions [Wang2008], and the behavior of mechanical systems [Chan2008, Dolleman2019]. A thorough understanding of activated escape from a metastable state provides insight into non-equilibrium phenomena that arise from the interplay between nonlinearities and fluctuations [Huber2020, McCann1999, Hanggi1986].
For this reason, numerous theoretical and experimental studies across various fields focus on computing and measuring the energy landscape of a system. The escape dynamics of a silica nanoparticle trapped in a bistable potential is extracted from the decaying autocorrelation of the which-well measurement [Militaru2021]. However, the measurement itself is not inherently well-suited to infer the experimental potential, so models are employed alongside the data to accurately replicate the landscape across all possible configurations, aiming at predicting the turnover. The (quasi-)energy surface of a micro-electromechanical resonator is reconstructed by Dumont et al. through ringdown measurements, revealing attractors and separatrices of the vector field. A model that accurately describes the underlying physics is then retrieved in the form of a Hamiltonian [Dumont2024]. Mello and Barrick experimentally obtained the protein energy landscape by measuring the stabilities of folded fragments of a repeat protein. The study of the energy surface as a function of the degree of folding of various parts of the protein reveals an early free energy barrier and suggests preferred low-energy pathways for folding [Mello2004]. A 1D Ising model is then employed to explore unfolding transitions. The landscape of a single unmodified protein, quantifying its change with temperature, is showcased in Peters et al. [Peters2024]. Kramers’ theory of transition rates is used to model the dynamics, demonstrating strong agreement with the observed state transitions. The potential energy landscape is able to capture the complex phenomenology of disordered materials. An energy evolution model has been introduced to describe the dynamics of the inherent structure in glassy systems [Fan2017]. Evaluating the energy landscape is crucial for understanding the presence of nonlinear processes that can give rise to critical transitions or tipping points [Ashwin2012, ArmstrongMcKay2022], which occur in various subsystems (tipping elements) of the Earth’s climate system [Chapman2025]. This brief selection of examples, while not exhaustive (cf. [Schon2024, Suzuki2021, Sanchez2024]), highlights the critical importance of accurate modeling in conjunction with precise measurements. Therefore, substantial effort is dedicated to obtaining a mathematical description of the landscape, which serves as a natural metric for assessing stability [Manuel2015, Gupta2011].
@@
In specific scenarios, the task of determining the representation of the energy landscape coincides with identifying the potential function. In gradient systems, the energy landscape directly corresponds to the potential function, meaning that the features and variations of the landscape, such as wells, barriers, and saddle points, are determined by the values and gradients of the potential function. However, when attempting to model the non-equilibrium statistics of transitions between stable states, it is common to encounter non-gradient systems [Cameron2012]. In this case, the transition dynamics becomes predictable and is entirely defined by the quasipotential that extends the idea of energy functions [Freidlin1998]. As a result, the quasipotential landscape provides an intuitive picture of the fundamental dynamical characteristics of complex systems operating outside of equilibrium [Dykman1998]. With the quasipotential at hand, valuable asymptotic information can be easily estimated. For instance, a calculation over the energy surface allows for the straightforward determination of the maximum likelihood transition path [Kim2015]. Additionally, a precise estimate for the invariant probability distribution near an attractor and the expected escape time from a basin of attraction has been formulated on the basis of the quasipotential topology [Bouchet2016]. The quasipotential holds considerable importance as an informative landscape in stochastic dynamical systems, providing insights into the system’s equilibria, the transition pathways between metastable states, and the rates of these transitions.
When trying to compute the quasipotential, one encounters a functional minimization problem that can be solved analytically only in specific special cases. In [Cameron2012], Cameron proposed a quasi-potential solver based on the ordered upwind method [Sethian2001, Sethian2004], which was inspired by Dijkstra’s algorithm [Sethian1996, Potter2019]. A significant development of the ordered upwind method for solving the Hamilton–Jacobi equation led to a new family of methods for computing the quasi-potential on regular meshes, known as the ordered line integral methods (OLIM) [Dahiya2018]. An extension and upgrade of OLIM, incorporating the midpoint quadrature rule, is the olim3D solver [Yang2019], which efficiently computes the quasi-potential on regular rectangular meshes in 3D. The minimum action methods have also been developed for computing the quasipotential between two specified points by solving a constrained minimization problem of the Freidlin-Wentzell action functional over the path space [Weinan2004, Zhou2008]. A Ritz discretization is used to minimize the action functional in [Kikuchi2020]. The geometric minimum action method [Heymann2020] provides an efficient way for solving the minimization problem based on the arclength parameterization of the path.
Although these approaches have their merits and numerous applications, we propose an alternative solution that does not require solving the variational formulation based on the Freidlin-Wentzell action functional or the associated Hamilton-Jacobi equation over a spatial mesh. Specifically, we employ data-driven methods and machine learning-based approaches for this purpose. The emerging scaling of these techniques has significantly impacted the field of nonlinear dynamics, showcasing their potential for both interpreting complex systems and developing innovative identification methods [Tang2020, Szalai2023, Chandrashekar2022]. In this study, we introduce a combination of data-driven approaches to achieve the identification and a symbolic interpretation of the quasipotential from data.
@@
To map the quasipotential landscape from data, a computational method based on neural networks was proposed in [lin2021data]. The method identifies an orthogonal decomposition of the vector field that governs the dynamics, in which the quasipotential can be inferred from the potential component. The idea of utilizing a decomposition of the vector field was also used to compute the generalized potential or invariant distribution for randomly perturbed dynamical systems [lin2022computing, lin2023computing]. In [Li2022], a neural network with automatic differentiation inspired by physics-informed neural network is employed to compute the quasipotential based on the Hamilton-Jacobi equation. Estimating the quasipotential with neural networks provides a powerful tool for strategically calculating the mean exit time in a broad class of stochastic dynamical systems [Li2023].
Here, we leverage the approach in [lin2021data] to compute a neural network solution of the quasipotential but improve the approach by incorporating normalization and rescaling steps into the input and output layers of the neural networks, respectively. The neural network solution serves as a preparatory step before integrating a constrained regression approach, which can reinterpret the black-box model by neural networks and yield an analytical model. The quasipotential equation adopts a parsimonious and interpretable form, meaning the model is sparse within the space of all possible functions [Kaptanoglu2021, Champion2020], without being restricted to any specific one. This sparse regression technique builds on the frame of the sparse identification of nonlinear dynamics (SINDy) introduced by [Brunton2016]. However, we adapt the optimization process to address the specific challenges of the quasipotential problem, particularly focusing on the identification of an involved target matrix. Our routine includes adjustments to preserve orthogonality while ensuring consistency with the original data. Notably, the combined method requires only observed trajectories of the dynamical system as input. Thus, the primary limitation of the technique lies in the availability of such trajectories, which may come directly from observations. By employing a fully data-driven approach, we derive a symbolic quasipotential function without imposing any predefined model for the energy landscape. The scalability of neural networks and sparse regression ensures that our approach is not limited by system dimensionality, making it readily extensible to high-dimensional systems.
The paper is organized as follows. In Sec. 2, we introduce the background of the quasipotential for non-gradient systems and the two parts of the data-driven method. In Sec. 3, we present two applications of the method. We also test it in the context of noisy dynamics, mimicking a realistic imperfect dataset, and extend its applicability to a four-dimensional system. In Sec. 4, we draw the conclusions. The implementation of the method, along with the presented examples, is available under the MIT License in the GitHub repository https://github.com/LinBoNUS/SIQ.
@@ 2.1 Expanding on the concept of potential function
We undergo the investigation and interpretation of the process through which a dynamical system evolves in presence of multiple equilibria. The underlying physics of the transient dynamics and attractors is assumed to be described by the state equation
@@
$\mathbf{f}(\mathbf{x})=-\nabla U(\mathbf{x}),$ (2)
we are in the presence of a gradient system, with $U$ referred to as the “potential”. Thanks to this function, the dynamics of this system can be likened to that of a ball-in-cup setup, where the surface is defined by the potential. In a more general scenario, i.e. non-gradient systems, it is not possible to find $U$ such that Eq. (2) is satisfied. However, the concept of a potential landscape can be extended to non-gradient systems through the Freidlin–Wentzell quasipotential function [Freidlin1998], which plays an analogous role in describing the system’s long-term stochastic behavior. Freidlin–Wentzell theory provides a framework for understanding rare noise-driven transitions between metastable states in systems perturbed by small random noise. Central to this theory is the action functional, which quantifies the likelihood of a stochastic trajectory over a given time interval. Let $\mathbf{x}_{0}$ be an attractor of the system (1), the quasipotential with respect to $\mathbf{x}_{0}$ is defined as
$U(\mathbf{x})=\inf_{T>0}\inf_{\varphi(t)}\int_{0}^{T}|\dot{\varphi}(t)-\mathbf{f}(\varphi(t))|^{2}dt,$ (3)
where the infimum is over all time $T>0$ and absolutely continuous path $\varphi(t)$, $0\leq t\leq T$ connecting $\mathbf{x}_{0}$ and $\mathbf{x}$. An alternative characterization for the quasipotential can be derived from a decomposition of the vector field [Brackston2018],
$\mathbf{f}(\mathbf{x})=-\nabla V(\mathbf{x})+\mathbf{g}(\mathbf{x}),$ (4)
@@
Assume that the function $V$ is continuously differentiable in $\mathcal{D}\cup\partial\mathcal{D}$ where $\mathcal{D}$ is a bounded domain in $\mathbb{R}^{d}$ and attains its strict local minimum at a point $A\in\mathcal{D}$. If a further condition that
[…]
holds, then the function $V$ is a scalar multiple of the quasipotential $U$ with respect to $A$ in a local region containing $A$ [Freidlin1998, lin2021data]. Specifically, that is $U(\mathbf{x})=2V(\mathbf{x})+C$ in the region $\{\mathbf{x}\in\mathcal{D}\cup\partial\mathcal{D}:V(\mathbf{x})\leq\min_{\mathbf{y}\in\partial\mathcal{D}}V(\mathbf{y})\}$, where $C$ is a constant.
The quasipotential represents a useful generalization of the potential function for a non-gradient system. In terms of the vector field, the $\nabla V$ component in Eq. (4) spurs a ball to roll to the bottom of a valley in the ball-in-a-cup analogy, thus $V$ specifies a hyper-surface in which all the trajectories move “downhill” in the absence of perturbations. From Eq. (4) and the Hamilton-Jacobi relation (5) it is straightforward to verify that
@@
hence, $\mathbf{g}$ and $\nabla V$ are perpendicular. Without additional external forces, the “circulatory” component $\mathbf{g}$ creates the circulation of trajectory around levels of $V$. For convenience, we refer to $\nabla V(\mathbf{x})$ and $\mathbf{g}(\mathbf{x})$ in the decomposition (4), (7) as downhill and circulatory components, respectively [Cameron2012, Freidlin1998].
Our method derives $V$ without solving numerically Eq. (5). A two-step data-driven approach is designed to output the symbolic expression of the quasipotential in terms of the state space variables, i.e. $V(\mathbf{x})$. In the first part of the method, a neural network determines the elements in the orthogonal decomposition of Eq.(4) (see Sec.2.2). Neural networks are recognized as universal function approximators, capable of approximating any continuous function on a compact domain to arbitrary accuracy, given sufficient capacity and a suitable architecture [Goodfellow2016, alpaydin2020]. This property makes them particularly effective for learning complex, nonlinear relationships between inputs and outputs without requiring an explicit analytical model.
In the second part of the method, the learned mapping functions, which are not associated with a physically interpretable set of equations, are reinterpreted via constrained regression (see Sec. 2.3). This technique identifies concise dynamic models by expressing the vector field as linear combinations of a large library of candidate functions and promoting sparsity in the coefficients through methods such as sequential threshold least squares or the lasso. Popularized by the SINDy framework [Brunton2016], this approach automatically selects only the essential terms required to reproduce the observed time series, yielding interpretable models that minimize overfitting. By discarding negligible basis functions, sparse regression provides direct insight into the underlying governing equations and remains computationally efficient even for high-dimensional systems.
@@ 2.2 Dataset and neural networks
$L(\theta)=$ $\frac{1}{NM}\sum_{i,j}\bigg\lVert\frac{1}{h}\Big(\mathcal{I}_{h}(X_{i}^{j,0};\mathbf{f}_{\theta})-X_{i}^{j,h}\Big)\bigg\rVert_{2}^{2}+\frac{\lambda}{S}\sum_{k=1}^{S}w\left(\frac{\nabla V_{\theta}(\tilde{X}_{k})\cdot\mathbf{g}_{\theta}(\tilde{X}_{k})}{|\nabla V_{\theta}(\tilde{X}_{k})|\cdot|\mathbf{g}_{\theta}(\tilde{X}_{k})|},\delta\right).$ (9)
@@
In Eq. (9), the piecewise function $w(z,\delta)=z^{2}\mathbb{I}_{z\geq 0}+\delta z^{2}\mathbb{I}_{z<0}$ with $\delta=0.1$, $\lambda$ is a parameter controlling the strength of the loss term corresponding to the orthogonality condition, and $\{\tilde{X}_{k}\}$, $1\leq k\leq S$ is a representative subset of $\{X_{i}^{j,0}\}$ with a uniform distribution in the state space.
We parameterize the potential component as $V_{\theta}(\mathbf{x})$ by the sum of a vanilla neural network $V^{\text{NN}}_{\theta}(\mathbf{x})$ with hyperbolic tangent ($\tanh$) as the activation function and a harmonic function [lin2021data]. This parameterization ensures the properties that $V_{\theta}$ is a radially unbounded function, and the set $H=\left\{\mathbf{x}\in\mathbb{R}^{d}:\nabla V_{\theta}(x)=0\right\}$ is a bounded and Lebesgue measure-zero set in $\mathbb{R}^{d}$. As a subset of $H$, the set of all minimizers of the function $V_{\theta}(\mathbf{x})$ is also bounded and has the measure zero. Thus, $V_{\theta}(\mathbf{x})$ serves as a candidate function with good physical properties for approximating the quasipotential.
@@
In addition, the training performances with the loss function (9) are ameliorated by adding the following two steps in the initialization of the neural networks: i) a normalization step to the input layer of $V_{\theta}$, $\mathbf{g}_{\theta}$ with respect to the data points; ii) a scalar multiplication to the output of the networks. In details, we construct the two networks $V_{\theta}$ and $\mathbf{g}_{\theta}$ as
@@
$V_{\theta}(\mathbf{x})=\eta_{v}\cdot\tilde{V}_{\theta}\left(\frac{\mathbf{x}-\mu}{\sigma}\right),\quad\text{in which}\ \tilde{V}_{\theta}(\mathbf{y})=V^{\text{NN}}_{\theta}(\mathbf{y})+|\mathbf{y}|^{2}$ (10)
and
@@
$\mathbf{g}_{\theta}(\mathbf{x})=\eta_{g}\cdot\mathbf{g}^{\text{NN}}_{\theta}\left(\frac{\mathbf{x}-\mu}{\sigma}\right),$ (11)
with $\mathbf{g}_{\theta}^{\text{NN}}$ being a vanilla neural network mapping from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ and $\mu$, $\sigma$ being the sample mean and deviation of the data points $\{X_{i}^{j,0}\}$, respectively. The parameters $\eta_{v}$, $\eta_{g}$ rescale the initial magnitudes of $|\nabla V_{\theta}|$, $|\mathbf{g}_{\theta}|$ to that of $|\mathbf{f}|$ estimated from the data. This is obtained by
$\eta_{v}=\operatorname*{arg\,min}_{\eta>0}\sum_{i,j}\Big(\big|\eta\mathbf{v}_{i}^{j}\big|-\big|Y_{i}^{j}\big|\Big)^{2}=\sum_{i,j}\big|\mathbf{v}_{i}^{j}\big|\cdot\big|Y_{i}^{j}\big|\bigg/\sum_{i,j}\big|\mathbf{v}_{i}^{j}\big|^{2}$ (12)
and
$\eta_{g}=\operatorname*{arg\,min}_{\eta>0}\sum_{i,j}\Big(\big|\eta\mathbf{g}_{i}^{j}\big|-\big|Y_{i}^{j}\big|\Big)^{2}=\sum_{i,j}\big|\mathbf{g}_{i}^{j}\big|\cdot\big|Y_{i}^{j}\big|\bigg/\sum_{i,j}\big|\mathbf{g}_{i}^{j}\big|^{2}.$ (13)
Here, $\mathbf{v}_{i}^{j}=\nabla\tilde{V}_{\theta}\big((X_{i}^{j,0}-\mu)/\sigma\big)$, $\mathbf{g}_{i}^{j}=\mathbf{g}^{\text{NN}}_{\theta}\big((X_{i}^{j,0}-\mu)/\sigma\big)$ and $Y^{j}_{i}=\big(X_{i}^{j,h}-X_{i}^{j,0}\big)/h$ is the estimated vector for $\mathbf{f}$ at the state $X_{i}^{j,0}$. Note that the parameters $\mu$, $\sigma$, $\eta_{v}$ and $\eta_{g}$ are kept as constant during the subsequent training process. This procedure has not been implemented in the earlier work [lin2021data].